OpenStudy (anonymous):

would 23√23/23 be √23?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

You can cancel all factors that the numerator and denominator have in common.

OpenStudy (anonymous):

can you multiply \[\sqrt{11}\] by \[\sqrt{3}\]

OpenStudy (anonymous):

Certainly: \[\sqrt{a} \cdot \sqrt{b} = \sqrt{a\cdot b}\]

OpenStudy (anonymous):

could i then divide \[\sqrt{33}\] by 3 ? or no because theres no \[\sqrt{?}\]

OpenStudy (anonymous):

Correct. \[\frac{\sqrt{33}}{3} \ne \sqrt{\frac{33}{3}}\]

OpenStudy (anonymous):

You cannot do that division.

OpenStudy (anonymous):

do you mean 23times by the square root of (23/23)? b/c 23/23 is just 1, so then the answer would be 23, not root 23

OpenStudy (anonymous):

YES